OpenStudy (anonymous):

How do I simplify d+7 / d^2-49 ????

OpenStudy (anonymous):

hint a^2-b^2=(a-b)(a+b)

OpenStudy (anonymous):

so d^2-49=(d-7)(d+7)

OpenStudy (anonymous):

@Eridanyo try to solve it and i will check ur answer

OpenStudy (anonymous):

@Abdulhameed okay thank you very much! I shall try it now

OpenStudy (anonymous):

do u want me to give u the answer ? so u can refer to it

OpenStudy (anonymous):

@Abdulhameed no I have to try to get it myself. I'm in honours math and I missed a day so I have to try to catch up tonight.

OpenStudy (anonymous):

@Abdulhameed I just needed the formula to help me get started.

OpenStudy (anonymous):

ok , good luck

OpenStudy (anonymous):

@Abdulhameed thank you.

OpenStudy (anonymous):

u r welcome , a medal plz lol

OpenStudy (anonymous):

thanx @Eridanyo